OpenStudy (anonymous):

Using graphing technology, approximate where the graph of y = 5x2 − 3x − 9 crosses the x-axis. (1.55, 0) and (−1.15, 0) (1.23, 0) and (−1.26, 0) (1.32, 0) and (−1.01, 0) (1.68, 0) and (−1.08, 0)

OpenStudy (anonymous):

Quadratic formula... a=5 b=-3 c=-9 Plug them in and you'll get two answers for x. Both y-values will be 0.

OpenStudy (anonymous):

Or do you want me to tell you how to do this on a TI-83/84?

OpenStudy (anonymous):

I have no idea how to do any of these..:/

OpenStudy (anonymous):

Do you have a graphing calculator?

OpenStudy (anonymous):

Nope...

OpenStudy (agreene):

the link I put is a picture of the graph--you can use it to do the estimation they are asking for.

OpenStudy (anonymous):

a=5 b=-3 c=-9 Plug them in and you'll get two answers for x. Both y-values will be 0.

OpenStudy (anonymous):

(-(-3) +/- sqrt((-3)^2 - 4*5*(-9)))/(2*5)

OpenStudy (anonymous):

(3 +/- (sqrt(9+180))/10 (3+/- sqrt(189))/10 ***THIS MEANS*** x = (3+ sqrt(189))/10 y=0 x = (3-sqrt(189))/10 y=0 Is this making sense?